OpenStudy (anonymous):

Graph each quadratic function y= x^2 -x...( what are the points to graph)

OpenStudy (espex):

Try plugging in some values for x. x | y ----- -3| -2| -1| 0| 1| 2| 3| Then record the corresponding y and plot the points.

OpenStudy (anonymous):

(-3,-8) would that be the first one

OpenStudy (espex):

Well the problem is y=x^2-x, so y = -3^2 -(-3) -> y = 9 + 3 -> y = 12 So the first point would be (-3,12)

OpenStudy (anonymous):

oh ok and the second would be (-2,6)

OpenStudy (anonymous):

and the third would be (-1,2)

OpenStudy (espex):

Looks right to me, yes. :)

OpenStudy (anonymous):

ok thank you

OpenStudy (espex):

You're welcome.